OpenStudy (gabylovesyou):

Which of the following is the slope between the points (2, 3) and (-3, 0)?

OpenStudy (gabylovesyou):

-5/3 5/3 -3/5 3/5

OpenStudy (anonymous):

\[ m = \frac{y_2 - y_1}{x_2 - x_1} = \frac{0 - 3}{-3 - 2} = \frac{-3}{-5} = \frac{3}{5} = 0.6 \]

OpenStudy (gabylovesyou):

so its d

OpenStudy (gabylovesyou):

thank you !!